Two buildings A and B adjacent to each other are separated by a narrow alley. Building A has a window at a heights of 12 meters and building B has a winder at a height of 8 meters, both windows facing the alley.
A 13meter ladder placed in the alley leaning on the wall of building A reaches the winder at a height of 12meters from the ground.
Without changing the position of the base of the ladder, the ladder is tilted towards building B and is able to just reach the winder at 8meters.
What is the width of the alley?

Answer by mananth(16949)

let the foot of the ladder be located at M
let the distance on ground from ladder foot to base of building A be MA
let the distance on ground from ladder foot to base of building B be MB
MA^2= 13^2-12^2
MA^2= 169-144=25
MA=5 m
...
MB^2= 13^2-8^2
MB^2=169-64
MB^2= 105
MB= 10.25 m
MA + MB = alley width
5+10.25= 15.25 m

m.ananth@hotmail.ca